# Any-Order Flexible Length Masked Diffusion

### v1

@@ Appendix D Joint Discrete Stochastic Interpolants: Definitions and Propositions for Section 3
Building on the discrete stochastic interpolant, we proceed to construct a masked diffusion model which natively supports growing in length by inserting mask tokens in arbitrary positions.
On a high-level, we would like define an interpolant constructed by deleting and masking sentences from the data distribution. However, the corresponding interpolating rate becomes hard to define, as it is no longer clear what each mask token should unmask to.
To this end, we introduce the joint interpolant mechanism that allows us to construct a broader class of interpolant and interpolating rate matrix by augmenting the interpolant with auxiliary information that allows us to specify more flexible interpolation path. We then leverage this newfound freedom to construct the flexible-length masked diffusion model.
@@ D.1 Joint Interpolant
By introducing an auxiliary variable coupled with the interpolant, joint interpolant expands the class of interpolating rate that can be defined.
@@
We note that this result is equivalent to ”Flow Matching with Auxiliary Process” in concurrent work Havasi et al. [2025].
@@ D.2 Flexible-Length Masked Diffusion
We then instantiate the joint interpolant to obtain the length-aware masked diffusion model, using a sorted list of indices that has been inserrted. Again, we drop $x_{0}$ in the writing as the model interpolates between a point mass at an empty sentence to the full data distribution.
@@
Let $x_{1}=(x_{1}^{0},\dots,x_{1}^{n-1})\sim p_{1}$ be a sequence of length $n$. Let $\alpha_{t}$ and $\beta_{t}$ be differentiable schedules on $[0,1]$ such that $\alpha_{0}=\beta_{0}=0$ and $\alpha_{1}=\beta_{1}=1$. Define insertion and unmasking times $\{T_{1}^{i}\}_{i=1}^{n}$, $\{T_{2}^{i}\}_{i=1}^{n}$ as follows:
@@
$s_{t}=\{i\in\{1,\dots,n\}\mid t>T_{1}^{i}\},$ (18)
with ascending order $s_{t}[0]<\dots<s_{t}[|s_{t}|-1]$, and boundary values:
$s_{t}[-1]=0,\quad s_{t}[|s_{t}|]=n,$ (19)
@@
Here, $s_{t}$ tracks the ordered set of indices whose tokens have been inserted and $s_{t}[i]$ is the $i$-th smallest element in $s_{t}$. The interpolant $x_{t}$ reveals the true token $x^{i}_{1}$ only after both insertion and unmasking.
Given access to this ordered set, one interpolating rate is:
Proposition D.2 (Length-Aware Masked Diffusion Interpolating Rate).
@@
    Unmask: For index set $s$, $x\subseteq x_{1}[s]$, $x^{i}=\mathbf{m}$, and $v\in\Sigma$:
@@
    Insert: For index set $s$, $x\subseteq x_{1}[s]$, $j\notin s$, and position $i$ such that s[i1]<j<s[i]s[i{-}1]<j<s[i]:
@@
$\partial_{t}p_{t}(s,x\mid x_{1})$ $=-(n-|s|)\dot{\alpha}_{t}(1-\alpha_{t})^{\,n-|s|-1}\prod_{i\in s}I_{i}(t)$
@@
Notice that this is equivalent to the R.H.S of the KFE (Eq. 3) if one uses the rate matrix $Q_{t}^{x_{1}}$. The four terms correspond to 1) Outgoing mass from insertion; 2) Outgoing mass from unmasking; 3) Incoming mass from unmasking; 4) Incoming mass from insertion. This concludes the proof.
@@
Parametrizing a approximate target rate matrix $\hat{R}_{t}$ in terms of an approximate per-token posterior $f_{\theta}(x,t)[i,v]\approx\mathbb{P}(x_{1}^{s_{t}[i]}=v\mid x_{t}=x)$, and an approximate number of insertion $g_{\theta}(x,t)[i]\approx\mathbb{E}_{s_{t}}\left[s_{t}[i]-s_{t}[i-1]-1\mid x_{t}=x\right]$. The target rate matrix can be learnt by minimising the following variational objective.
@@
The loss functional defined as:
$L[\hat{R}_{t}]$ $=\int_{0}^{1}\mathbb{E}_{x_{1},s_{t},x_{t}}\left[-\frac{\dot{\beta}_{t}}{1-\beta_{t}}\sum_{i=1}^{|x_{1}|}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,v]\right]\;dt,$ (25)
@@
Let $\mathbb{P}$ and $\mathbb$$\hat${P}} be the path measure of a continuous time markov chain starting with the empty string with rate matrix $R_{t}$ and $\hat{R_{t}}$ respectively.
@@
$=\int_{0}^{1}\mathbb{E}_{x_{1},s_{t},x_{t}}\left[-\frac{\dot{\beta}_{t}}{1-\beta_{t}}\sum_{i=1}^{|x_{1}|}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,v]\right]\;dt$
@@
The terminal KL-bound then follows from the data processing inequality that is:
@@ E.1 Precise detail on the inference algorithms
In this section, we provide details on the unmasking steps of vanilla and adaptive inference for FlexMDM, summarized in Algorithm 4. Suppose at inference time we are given the discretization step size $\tau$, a partially observed sequence $x_{t_{k}}$, and the current time step $t_{k}$.
@@
We first draw the number of tokens to unmask, denoted by an integer $K$. While Proposition 3 and Theorem E.1 show that the choice of $K$ does not affect the theoretical guarantees, in practice, we set $K$ to match the expected number of unmasked tokens under vanilla inference yields stable behavior. Accordingly, we sample $K\sim\mathrm{Poi}\!\left(\tau\cdot\tfrac{\dot{\beta}_{t_{k}}}{1-\beta_{t_{k}}}\cdot\#\{\text{masked tokens in }x_{t_{k}}\}\right)$.
@@
E.2 Proof of FlexMDM’s any-order inference capability.
@@ E.2.1 Proof preliminaries
Recall that in the definition of the joint interpolant we defined an index-tracking variable $s_{t}$ which essentially tracked which indices of $x_{1}$ correspond to the tokens in $x_{t}$. While our analysis below will not use the language of stochastic interpolants, we will still use the idea of tracking $s_{t}$, with slightly modified notation. Specifically, for any $0\leq t<1$, we will use the notation $\Pr_{(x_{1},s)\mid x_{t}=x}$ and $\mathbb{E}_{(x_{1},s)\mid x_{t}=x}$ to denote probability and expectation with respect to the distribution given by sampling $x_{1}$ from $q_{t}(\cdot\mid x)$, and then sampling $s$ uniformly random from subsets for which $x\subseteq(x_{1})|_{s}$. When we only care about the marginal distribution over $s$, we write $\Pr_{s\mid x_{t}=x}$ and $\mathbb{E}_{s\mid x_{t}=x}$. Given such a subset $s$ and $i\in\{1,\ldots,|s|\}$, we use $s_{i}$ to denote its $i$-th element in sorted order. The insertion expectations which we had denoted by $\mathbb{E}[s_{t}[i]-s_{t}[i-1]-1]$ in the main body are thus given by $\mathbb{E}_{s\mid x_{t}=x}[s_{i}-s_{i-1}-1]$ in the notation of this section.
@@ E.3 Formal guarantee for adaptive inference
    $R^{\rm ins}_{t}(x,y)=\begin{cases}\mathbb{E}_{s\mid x_{t}=x}[s_{i+1}-s_{i}-1]\cdot\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}&\text{if}\ y=x\,\triangleleft_{i}\mathbf{m}\\ -\sum^{\mathrm{len}(x)}_{i=0}R^{\rm ins}_{t}(x,x\,\triangleleft_{i}\mathbf{m})&\text{if}\ y=x\\ 0&\text{otherwise}\end{cases}$ (29)
@@
$\frac{\sum_{x_{1}}p(x_{1})\cdot(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i+1}-s_{i}-1)}{\sum_{x_{1}}p(x_{1})\cdot(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\#\{S:x\subseteq(x_{1})|_{S}\}}\cdot\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\,.$ (33)
Furthermore, $\sum_{i}(s_{i+1}-s_{i}-1)=\mathrm{len}(x_{1})-\mathrm{len}(x)$, so
@@
$\sum^{\mathrm{len}(x)}_{i=0}\frac{\sum_{x_{1}}p(x_{1})(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i+1}-s_{i}-1)}{\sum_{x_{1}}p(x_{1})(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\#\{S:x_{t}\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S}\}}\cdot\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq x^{*}|_{S}\}\\ =\#\{S:x\subseteq(x^{*})|_{S}\}\cdot(\mathrm{len}(x^{*})-\mathrm{len}(x))$ (37)
The key combinatorial step is as follows. For any $x_{1}$ in the support of $p$, consider a subset $S$ for which $x\subseteq(x_{1})|_{S}$. Note that for every such $S$, we can uniquely associate exactly $s_{i+1}-s_{i}-1$ different subsets $S^{\prime}$ of size $|S|+1$ for which $x\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S^{\prime}}$. Therefore, $\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i+1}-s_{i}-1)=\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S}\}$, and the left-hand side of Eq. Eq. (37) thus becomes
$\sum^{\mathrm{len}(x)}_{i=0}\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq x^{*}|_{S}\}=\sum^{\mathrm{len}(x)}_{i=0}\sum_{S:x\subseteq x^{*}|_{S}}(s_{i+1}-s_{i}-1)=\sum_{S:x\subseteq(x_{1})|_{S}}(\mathrm{len}(x^{*})-\mathrm{len}(x_{t}))\,$, (38)
@@ F.1 Pretraining on OpenWebText
For evaluation, we take sequences generated by both models and retain the clean tokens by removing padding (e.g., the leading pad token). We adopt LLaMA-2-7B [Touvron et al., 2023] as the reference model to compute likelihoods. We notice that MDM struggles to generate short sentences with low perplexities. Therefore, we filter overly short sequences with $\leq$ 10 tokens when calculating average perplexity.
@@ F.2 Pretraining on the Maze planning task
We generate mazes with a fully random, recursive division procedure (Listing 1), on a $41\times 41$ grid where some cells are invalid. As described in Section 5.1, we use a subgoal-conditioned planning task: the model is given a sequence of subgoals $(g_{1},\dots,g_{K})$ and must produce a valid path that connects them in order. To construct training examples for a given $K$, we sample $15000$ start–goal pairs, compute the shortest path for each pair via breadth-first search, and then add controlled detours to obtain up to $10$ distinct valid paths per pair. Subgoals are formed by selecting K2K-2 intermediate cells uniformly at random along a chosen path (start and goal are already fixed). We use $95\%$ of the pairs for training and hold out $5\%$ for validation to evaluate generalization to unseen pairs and subgoal sets.
@@
For MDM, the training sequence is $((g_{1},\dots,g_{K})\;\texttt{[SEP]}\;\texttt{Path})$, where [SEP] is a special separator and Path denotes the tokenized path. During training, the prompt $(g_{1},\dots,g_{K})$ bypasses the interpolant so that, at inference time, the model can condition on $(g_{1},\dots,g_{K})\;\texttt{[SEP]}$ and generate the path. For FlexMDM, we use an interpolant in which the subgoal tokens are exempt from the process in Eq. (6); that is, tokens corresponding to each $g_{i}$ are kept clean at all times. This also enables generation to start from $(g_{1},\dots,g_{K})$.
@@
We use the same design as in the OpenWebText pretraining but with a smaller model: hidden size:256, heads:8, conditional dimension:128, dropout:0.1, layers:8. Both models have approximately 8.58.5M parameters. We train on $4$ A100 GPUs with a global batch size of $256$ for up to $100$ epochs. We use AdamW [Loshchilov and Hutter, 2017] with weight decay $0.01$, learning rate $1\times 10^{-4}$, $20$ warmup epochs, and an EMA factor of $0.9999$.
@@
Given the final conditionally generated sequence, we report success rate under two criteria: (1) all visited cells are valid (no collisions with blocked cells), and (2) the path connects the subgoals consecutively in order.

### v2

@@ Appendix D Joint Discrete Stochastic Interpolants: Definitions and Propositions for Section 3
Building on the discrete stochastic interpolant, we proceed to construct a discrete diffusion that models a probability distribution whose supports span variable-length sequences.
On a high level, we would like to define an interpolant constructed by deleting and masking sentences from the data distribution. However, the corresponding interpolating rate becomes cumbersome to characterize, as it is no longer clear what each mask token should unmask to.
To this end, we introduce the joint interpolant that allows us to construct a broader class of interpolants and interpolating rate matrices by augmenting the interpolant with auxiliary information that allows us to specify a more flexible interpolation path. We then leverage this newfound freedom to construct the flexible-length masked diffusion model.
@@ D.1 Joint Interpolant
By introducing an auxiliary variable coupled with the interpolant, the joint interpolant expands the class of interpolating rates that can be defined.
@@
We note that this result is equivalent to “Flow Matching with Auxiliary Process” in concurrent work Havasi et al. [2025].
@@ D.2 Flexible-Length Masked Diffusion
We then instantiate the joint interpolant to obtain the length-aware masked diffusion model, using a sorted list of indices that has been inserted. Again, we drop $x_{0}$ in the writing as the model interpolates between a point mass at an empty sentence to the full data distribution. We redefine the interpolant in equation 6 for clarity.
@@
Let $x_{1}=(x_{1}^{0},\dots,x_{1}^{n-1})\sim p_{1}$ be a sequence of length $n$. Let $\alpha_{t}$ and $\beta_{t}$ be monotone and differentiable schedules on $[0,1]$ such that $\alpha_{0}=\beta_{0}=0$ and $\alpha_{1}=\beta_{1}=1$. Define insertion and unmasking times $\{T_{1}^{i}\}_{i=0}^{n-1}$, $\{T_{2}^{i}\}_{i=0}^{n-1}$ as follows:
@@
$s_{t}=\{i\in\{0,\dots,n-1\}\mid t>T_{1}^{i}\},$ (18)
with ascending order $s_{t}[0]<\dots<s_{t}[\mathrm{len}(s_{t})-1]$, and boundary values:
$s_{t}[-1]=-1,\quad s_{t}[\mathrm{len}(s_{t})]=n,$ (19)
@@
Here, $s_{t}$ tracks the ordered set of indices whose tokens have been inserted. The interpolant $x_{t}$ reveals the true token $x^{i}_{1}$ only after both insertion and unmasking. Given access to this ordered set, one interpolating rate is:
Proposition D.2 (FlexMDM Interpolating Rate).
@@
    Unmask: For index set $s$, $x\subseteq{x_{1}}|_{s}$, $x^{i}=\mathbf{m}$, and $v\in\Sigma$:
@@
    Insert: For index set $s$, $x\subseteq{x_{1}}|_{s}$, $j\notin s$, and position $i$ such that s[i1]<j<s[i]s[i{-}1]<j<s[i]:
@@
$\partial_{t}p_{t}(s,x\mid x_{1})$
$=-(n-|s|)\dot{\alpha}_{t}(1-\alpha_{t})^{\,n-|s|-1}\prod_{i\in s}I_{i}(t)$
@@
Notice that this is equivalent to the R.H.S of the KFE (Eq. 3) if one uses the rate matrix $Q_{t}^{x_{1}}$. The four terms correspond to 1) Outgoing mass from insertion; 2) Outgoing mass from unmasking; 3) Incoming mass from unmasking; 4) Incoming mass from insertion. This concludes the proof. ∎
@@
Performing an approximate target rate matrix $\hat{R}_{t}$ in terms of an approximate posterior by token $f_{\theta}(x,t)[i,v]\approx\mathbb{P}(x_{1}^{s_{t}[i]}=v\mid x_{t}=x)$ and an approximate number of insertions $g_{\theta}(x,t)[i]\approx\mathbb{E}_{s_{t}}\left[s_{t}[i]-s_{t}[i-1]-1\mid x_{t}=x\right]$. The target rate matrix can be learned by minimizing the following variational objective. Note that the variational loss objective below is the same as one we defined in equation 7.
@@
The loss function is defined as:
$L[\hat{R}_{t}]$ $=\int_{0}^{1}\mathbb{E}_{x_{1},s_{t},x_{t}}\left[-\frac{\dot{\beta}_{t}}{1-\beta_{t}}\sum_{i=0}^{\mathrm{len}(x_{t})-1}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{s_{t}[i]}]\right]\;dt,$ (25)
@@
Let $\mathbb{P}$ and $\mathbb$$\hat${P}} be the path measure of a continuous time Markov chain starting with the empty string with rate matrix $R_{t}$ and $\hat{R_{t}}$, respectively.
@@
$=\int_{0}^{1}\mathbb{E}_{x_{1},s_{t},x_{t}}\left[-\frac{\dot{\beta}_{t}}{1-\beta_{t}}\sum_{i=0}^{\mathrm{len}(x_{t})-1}\mathbf{1}\{x_{t}^{i}=\mathbf{m}\}\log f_{\theta}(x_{t},t)[i,x_{1}^{s_{t}[i]}]\right]\;dt$
@@
The terminal KL-bound then follows from the data processing inequality, that is:
@@ E.1 Precise detail on the inference algorithms
In this section, we provide details on the unmasking steps of vanilla and adaptive inference for FlexMDM, summarized in Algorithm 4, Subroutine 2. Suppose at inference time we are given the discretization step size $\tau$, a partially observed sequence $X_{t_{k}}$, and the current time step $t_{k}$.
@@
We first draw the number of tokens to unmask, denoted by an integer $K$. While Proposition 3 (Theorem E.1) shows that the choice of $K$ does not affect the theoretical guarantees, in practice, we set $K$ to match the expected number of unmasked tokens under vanilla inference yields stable behavior. Accordingly, we sample $K\sim\mathrm{Poi}\!\left(\tau\cdot\tfrac{\dot{\beta}_{t_{k}}}{1-\beta_{t_{k}}}\cdot\#\{\text{masked tokens in }X_{t_{k}}\}\right)$.
@@
E.2 Proof of FlexMDM’s any-order inference capability
@@ E.2.1 Proof preliminaries
Recall that in the definition of the joint interpolant we defined an index-tracking variable $s_{t}$ which essentially tracked which indices of $x_{1}$ correspond to the tokens in $x_{t}$. While our analysis below will not use the language of stochastic interpolants, we will still use the idea of tracking $s_{t}$, with slightly modified notation. Specifically, for any $0\leq t<1$, we will use the notation $\Pr_{(x_{1},s)\mid x_{t}=x}$ and $\mathbb{E}_{(x_{1},s)\mid x_{t}=x}$ to denote probability and expectation with respect to the distribution given by sampling $x_{1}$ from $q_{t}(\cdot\mid x)$, and then sampling $s$ uniformly random from subsets for which $x\subseteq(x_{1})|_{s}$. When we only care about the marginal distribution over $s$, we write $\Pr_{s\mid x_{t}=x}$ and $\mathbb{E}_{s\mid x_{t}=x}$. Given such a subset $s$ and $i\in\{0,\ldots,|s|-1\}$, we use $s_{i}$ to denote its $i$-th element in sorted order; as before, we also define the boundary values $s_{-1}=-1$ and $s_{\mathrm{len}(s)}=\mathrm{len}(x_{1})$. The insertion expectations which we had denoted by $\mathbb{E}[s_{t}[i]-s_{t}[i-1]-1]$ in the main body are thus given by $\mathbb{E}_{s\mid x_{t}=x}[s_{i}-s_{i-1}-1]$ in the notation of this section.
@@ E.3 Formal guarantee for adaptive inference
    $R^{\rm ins}_{t}(x,y)=\begin{cases}\mathbb{E}_{s\mid x_{t}=x}[s_{i}-s_{i-1}-1]\cdot\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}&\text{if}\ y=x\,\triangleleft_{i}\mathbf{m}\\ -\sum^{\mathrm{len}(x)}_{i=0}R^{\rm ins}_{t}(x,x\,\triangleleft_{i}\mathbf{m})&\text{if}\ y=x\\ 0&\text{otherwise}\end{cases}$ (29)
@@
$\frac{\sum_{x_{1}}p(x_{1})\cdot(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i}-s_{i-1}-1)}{\sum_{x_{1}}p(x_{1})\cdot(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\#\{S:x\subseteq(x_{1})|_{S}\}}\cdot\frac{\dot{\alpha}_{t}}{1-\alpha_{t}}\,.$ (33)
Furthermore, $\sum_{i}(s_{i}-s_{i-1}-1)=\mathrm{len}(x_{1})-\mathrm{len}(x)$, so
@@
$\sum^{\mathrm{len}(x)}_{i=0}\frac{\sum_{x_{1}}p(x_{1})(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i}-s_{i-1}-1)}{\sum_{x_{1}}p(x_{1})(1-\alpha_{t})^{\mathrm{len}(x_{1})-\mathrm{len}(x)}\cdot\#\{S:x_{t}\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S}\}}\cdot\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq x^{*}|_{S}\}\\ =\#\{S:x\subseteq(x^{*})|_{S}\}\cdot(\mathrm{len}(x^{*})-\mathrm{len}(x))$ (37)
The key combinatorial step is as follows. For any $x_{1}$ in the support of $p$, consider a subset $S$ for which $x\subseteq(x_{1})|_{S}$. Note that for every such $S$, we can uniquely associate exactly $s_{i}-s_{i-1}-1$ different subsets $S^{\prime}$ of size $|S|+1$ for which $x\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S^{\prime}}$. Therefore, $\sum_{S:x\subseteq(x_{1})|_{S}}(s_{i}-s_{i-1}-1)=\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq(x_{1})|_{S}\}$, and the left-hand side of Eq. Eq. (37) thus becomes
$\sum^{\mathrm{len}(x)}_{i=0}\#\{S:x\,\triangleleft_{i}\mathbf{m}\subseteq x^{*}|_{S}\}=\sum^{\mathrm{len}(x)}_{i=0}\sum_{S:x\subseteq x^{*}|_{S}}(s_{i}-s_{i-1}-1)=\sum_{S:x\subseteq(x_{1})|_{S}}(\mathrm{len}(x^{*})-\mathrm{len}(x_{t}))\,$, (38)
@@ F.1 Pretraining on OpenWebText
For evaluation, we take sequences generated by both models and retain the clean tokens by removing padding (e.g., the leading pad token). We adopt LLaMA-2-7B [Touvron et al., 2023] as the reference model to compute likelihoods. We notice that the pretrained MDM generates short sentences with unreasonably large (worse) perplexities. Therefore, we filter overly short sequences with $\leq$ 10 tokens when calculating average perplexity.
@@ F.2 Pretraining on the Maze planning task
We generate mazes with a fully random, recursive division procedure (the code is provided in Listing 1), on a $41\times 41$ grid, with some invalid cells. As described in Section 5.1, we consider a subgoal-conditioned planning task: the model is given a sequence of subgoals $(g_{1},\dots,g_{K})$ and must produce a valid path that connects them in order. To construct training examples for a given $K$, we sample $15000$ start–goal pairs ($g_{1},g_{K}$), compute the shortest path for each pair via breadth-first search, and then add controlled detours to obtain up to $10$ distinct valid paths per pair. Subgoals are formed by selecting K2K-2 intermediate cells uniformly at random along a chosen path (start and goal are already fixed). We use $95\%$ of the pairs for training and hold out $5\%$ for validation to evaluate generalization to unseen pairs and subgoal sets.
@@
For MDM, the training sequence is $((g_{1},\dots,g_{K})\;\texttt{[SEP]}\;\texttt{Path})$, where [SEP] is a special separator and Path denotes the tokenized path. During training, the prompt $(g_{1},\dots,g_{K})$ bypasses the interpolant so that, at inference time, the model can condition on $(g_{1},\dots,g_{K})\;\texttt{[SEP]}$ and generate the path. For FlexMDM, we use an interpolant in which the subgoal tokens are exempt from the process in Eq. (6); that is, tokens corresponding to each $g_{i}$ are kept clean at all times. This also enables generation to start from $(g_{1},\dots,g_{K})$. Although this conditional generation template changes the base distributions $p_{0}$ for both MDM and FlexMDM, we note that the theoretical guarantee from Section 3 and Section 4 still holds–once the training is perfect (under the access to the ground truth unmasking posterior and insertion expectation), the inference algorithms recover the ground truth distribution $p_{1}$.
@@
We use the same architectural design as in the OpenWebText pretraining, but with a smaller model: hidden size:256, heads:8, conditional dimension:128, dropout:0.1, layers:8. Both models have approximately 8.58.5M parameters. We train them on $4$ A100 GPUs with a global batch size of $256$ for up to $100$ epochs. We use AdamW [Loshchilov and Hutter, 2017] with weight decay $0.01$, learning rate $1\times 10^{-4}$, $20$ warmup epochs, and an EMA factor of $0.9999$.
@@
Given the final conditionally generated sequence, we report the success rate under two criteria: (1) all visited cells are valid (no collisions with invalid cells), and (2) the path connects the subgoals consecutively in order. We perform inference both models with the number of sampling steps $256$.